# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 2.4. Canonical correlation analysis
where $A$ and $B$ are non-degenerate $N\times N$ and $M\times M$ matrices, respectively, $I_{N}$ and $I_{M}$ are identity matrices of $N\times N$ and $M\times M$ dimensions, respectively, and $\mathrm{diag}(\sqrt{\theta_{1}},\dots,\sqrt{\theta_{r}})$ is a rectangular matrix with $\sqrt{\theta_{1}},\dots,\sqrt{\theta_{r}}$ on the first $r$ elements of the main diagonal and 00 everywhere else.
Let $\mathbf{x}$ be an $(N+M)$–dimensional Gaussian mean 00 random vector with covariance $\Omega$, and let $\mathbf{u}$ and $\mathbf{v}$ denote its first $N$ and last $M$ coordinates, respectively. The parameters $\theta_{1},\dots,\theta_{r}$ are the squared canonical correlations between $\mathbf{u}$ and $\mathbf{v}$; see Bykhovskaya and Gorin (2024), as well as classical statistics references such as Thompson (1984); Gittins (1985); Anderson (2003); Muirhead (2009) for detailed introductions to canonical correlation analysis (CCA). Algorithmically, $\theta_{i}$ are the largest eigenvalues of the matrix $(\mathbb{E}\mathbf{u}\mathbf{u}^{\mathsf{T}})^{-1}\mathbb{E}\mathbf{u}\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{T}})^{-1}\mathbb{E}\mathbf{v}\mathbf{u}^{\mathsf{T}}$.
Given $S$ independent samples of $\mathbf{x}$, we construct two matrices: the $N\times S$ matrix $\mathbf{U}$ has $S$ samples of $\mathbf{u}$ as its columns and the $M\times S$ matrix $\mathbf{V}$ has $S$ samples of $\mathbf{v}$ as its columns. The sample squared canonical correlations $\lambda_{1}\geq\lambda_{2}\geq\dots$ are the eigenvalues of the $N\times N$ matrix $(\mathbf{U}\mathbf{U}^{\mathsf{T}})^{-1}\mathbf{U}\mathbf{V}^{\mathsf{T}}(\mathbf{V}\mathbf{V}^{\mathsf{T}})^{-1}\mathbf{V}\mathbf{U}^{\mathsf{T}}$. Our goal is to estimate $\theta_{1},\dots,\theta_{r}$ from these observed eigenvalues.
@@ 3. Construction of confidence intervals
In this section we present our algorithm for constructing confidence intervals and explain how they can be interpreted and used to distinguish between noise, non-informative signals, and meaningful signals. We begin by introducing the transition process $\mathcal{T}(\Theta)$ and its properties, and then show how to use it to construct confidence intervals. The underlying theorems will be presented in Section 4.
@@ 3.1. Transition process
$\Theta$ $\alpha$005 .025 .05 .5 .95 .975 .995 -3.0 -3.85 -3.22 -2.89 -0.96 1.32 1.80 2.78 -2.9 -3.85 -3.22 -2.88 -0.95 1.32 1.80 2.79 -2.8 -3.84 -3.20 -2.86 -0.94 1.33 1.81 2.80 -2.7 -3.83 -3.20 -2.86 -0.93 1.35 1.83 2.82 -2.6 -3.82 -3.19 -2.85 -0.92 1.36 1.85 2.83 -2.5 -3.82 -3.18 -2.84 -0.91 1.38 1.86 2.83 -2.4 -3.80 -3.17 -2.83 -0.89 1.38 1.87 2.85 -2.3 -3.80 -3.16 -2.82 -0.89 1.42 1.91 2.90 -2.2 -3.79 -3.15 -2.82 -0.87 1.42 1.90 2.86 -2.1 -3.77 -3.13 -2.79 -0.85 1.44 1.93 2.94 -2.0 -3.75 -3.12 -2.78 -0.84 1.46 1.95 2.97 -1.9 -3.75 -3.11 -2.77 -0.82 1.49 1.98 2.98 -1.8 -3.74 -3.10 -2.75 -0.80 1.52 2.01 3.03 -1.7 -3.73 -3.09 -2.74 -0.78 1.54 2.03 3.05 -1.6 -3.71 -3.06 -2.72 -0.76 1.57 2.07 3.11 -1.5 -3.69 -3.05 -2.70 -0.74 1.61 2.11 3.14 -1.4 -3.67 -3.03 -2.69 -0.71 1.64 2.14 3.19 -1.3 -3.65 -3.01 -2.66 -0.69 1.68 2.19 3.26 -1.2 -3.64 -2.99 -2.65 -0.66 1.72 2.23 3.27 -1.1 -3.61 -2.97 -2.63 -0.63 1.77 2.29 3.37 -1.0 -3.58 -2.94 -2.59 -0.59 1.83 2.35 3.44 -0.9 -3.57 -2.91 -2.56 -0.56 1.88 2.41 3.52 -0.8 -3.55 -2.89 -2.54 -0.53 1.94 2.48 3.62 -0.7 -3.53 -2.87 -2.51 -0.48 2.01 2.57 3.70 -0.6 -3.51 -2.84 -2.49 -0.44 2.08 2.64 3.79 -0.5 -3.49 -2.82 -2.46 -0.39 2.17 2.74 3.93 -0.4 -3.45 -2.77 -2.42 -0.35 2.27 2.85 4.10 -0.3 -3.41 -2.74 -2.38 -0.28 2.38 2.97 4.25 -0.2 -3.37 -2.70 -2.33 -0.22 2.48 3.09 4.39 -0.1 -3.35 -2.66 -2.30 -0.15 2.61 3.24 4.59 0.0 -3.31 -2.62 -2.25 -0.08 2.75 3.40 4.76 0.1 -3.26 -2.57 -2.20 0.00 2.91 3.57 4.96 0.2 -3.22 -2.52 -2.14 0.09 3.08 3.77 5.24 0.3 -3.17 -2.47 -2.09 0.19 3.25 3.95 5.45 0.4 -3.13 -2.40 -2.02 0.29 3.46 4.19 5.72 0.5 -3.08 -2.34 -1.96 0.41 3.67 4.42 5.99 0.6 -3.01 -2.28 -1.89 0.54 3.90 4.68 6.29 0.7 -2.96 -2.21 -1.80 0.69 4.17 4.96 6.60 0.8 -2.90 -2.14 -1.73 0.83 4.41 5.23 6.91 0.9 -2.81 -2.04 -1.62 1.01 4.73 5.57 7.30 1.0 -2.71 -1.95 -1.53 1.18 5.02 5.88 7.62 1.1 -2.63 -1.86 -1.42 1.39 5.37 6.23 8.01 1.2 -2.56 -1.75 -1.31 1.60 5.69 6.59 8.38 1.3 -2.47 -1.64 -1.18 1.84 6.06 6.96 8.80 1.4 -2.37 -1.52 -1.05 2.10 6.44 7.37 9.22 1.5 -2.25 -1.38 -0.89 2.37 6.82 7.79 9.72 $\Theta$ $\alpha$005 .025 .05 .5 .95 .975 .995 1.5 -2.25 -1.38 -0.89 2.37 6.82 7.79 9.72 1.6 -2.14 -1.23 -0.73 2.67 7.27 8.22 10.19 1.7 -2.03 -1.08 -0.55 3.00 7.67 8.65 10.66 1.8 -1.88 -0.91 -0.37 3.34 8.12 9.13 11.19 1.9 -1.73 -0.72 -0.14 3.70 8.59 9.62 11.71 2.0 -1.56 -0.53 0.07 4.09 9.09 10.14 12.20 2.1 -1.39 -0.30 0.34 4.50 9.61 10.66 12.84 2.2 -1.21 -0.05 0.62 4.91 10.12 11.20 13.31 2.3 -0.99 0.22 0.92 5.37 10.67 11.75 13.96 2.4 -0.77 0.51 1.24 5.84 11.24 12.36 14.52 2.5 -0.53 0.82 1.59 6.32 11.82 12.94 15.14 2.6 -0.24 1.16 1.96 6.82 12.45 13.57 15.88 2.7 0.04 1.52 2.37 7.35 13.05 14.22 16.52 2.8 0.36 1.91 2.78 7.90 13.68 14.85 17.18 2.9 0.73 2.34 3.23 8.47 14.35 15.55 17.90 3.0 1.08 2.77 3.70 9.04 15.02 16.21 18.60 3.1 1.52 3.27 4.22 9.67 15.71 16.92 19.38 3.2 1.93 3.72 4.72 10.28 16.42 17.63 20.07 3.3 2.41 4.25 5.26 10.93 17.15 18.40 20.85 3.4 2.93 4.79 5.82 11.61 17.91 19.18 21.74 3.5 3.40 5.36 6.41 12.29 18.65 19.93 22.48 3.6 3.96 5.95 7.03 13.01 19.48 20.78 23.37 3.7 4.52 6.59 7.68 13.72 20.26 21.56 24.14 3.8 5.12 7.18 8.30 14.48 21.13 22.47 25.05 3.9 5.76 7.89 9.01 15.25 21.95 23.31 25.98 4.0 6.38 8.57 9.72 16.05 22.83 24.17 26.87 4.1 7.04 9.25 10.41 16.85 23.70 25.06 27.75 4.2 7.70 9.93 11.15 17.68 24.61 25.96 28.68 4.3 8.41 10.69 11.91 18.52 25.55 26.94 29.71 4.4 9.14 11.49 12.73 19.41 26.49 27.88 30.62 4.5 9.93 12.27 13.52 20.28 27.43 28.84 31.60 4.6 10.68 13.09 14.35 21.20 28.43 29.81 32.63 4.7 11.46 13.91 15.19 22.12 29.41 30.84 33.67 4.8 12.23 14.76 16.07 23.06 30.44 31.88 34.68 4.9 13.12 15.63 16.94 24.03 31.52 32.99 35.90 5.0 13.97 16.54 17.87 25.04 32.55 34.02 36.98 5.1 14.88 17.44 18.80 26.04 33.64 35.13 38.05 5.2 15.76 18.39 19.73 27.06 34.74 36.22 39.18 5.3 16.70 19.36 20.71 28.11 35.83 37.34 40.37 5.4 17.67 20.35 21.75 29.19 37.01 38.52 41.57 5.5 18.57 21.35 22.75 30.28 38.14 39.70 42.71 5.6 19.63 22.36 23.78 31.40 39.34 40.91 43.98 5.7 20.67 23.43 24.85 32.54 40.53 42.08 45.15 5.8 21.65 24.46 25.91 33.67 41.72 43.31 46.43 5.9 22.78 25.52 27.00 34.83 42.97 44.54 47.70 6.0 23.82 26.63 28.11 36.03 44.25 45.86 48.95
@@
Table 1. Quantiles of $\mathcal{T}(\Theta)$ for $-3\leq\Theta\leq 6$ based on $MC=10^{6}$ Monte Carlo simulations.
$\alpha$005 .025 .05 .5 .95 .975 .995
quantile $F_{1}^{-1}(\alpha)$4.15 -3.52 -3.18 -1.27 0.98 1.45 2.42
@@
  • For $-3\leq\Theta\leq 6$, quantiles are tabulated in Table 1 using the algorithm described in Section 4.1.
@@
    $\mathbb{P}\left\{\mathcal{T}(\Theta)\leq t\right\}\approx\Phi((t-\Theta^{2})/(2\sqrt{\Theta}))\leavevmode\nobreak\ .$
@@
    $\mathbb{P}\left\{\mathcal{T}(\Theta)\leq t\right\}\approx F_{1}(t+1/\Theta)\leavevmode\nobreak\ ,$
@@
Theorem 4.6 and Corollary 4.8 establish that the approximation (3.1) is valid both when $\theta$ is bounded away from the critical value $\theta^{c}$ and when $\theta$ is close to $\theta^{c}$. These results also show that, in many cases, the approximations for different $\lambda_{i}$ are asymptotically independent. Hence, we can use (3.1) as a foundation for constructing confidence intervals.
3.2. Confidence intervals with known $\sigma^{2}$
@@
The first step is to draw a histogram of all eigenvalues $\lambda_{1},\lambda_{2},\dots$. In the settings of (2.1), (2.6), (2.10), or (2.14), the histogram should closely resemble a known limiting shape; namely, the semicircle law, Marchenko-Pastur law, or Wachter law, depending on the model, as detailed in Table 3, with parameters specified in Table 4, see Appendix 8.4 for more details. If the histogram is reminiscent of one of these shapes, we regard the modelling assumptions as valid and apply Procedure 3.1 to construct confidence intervals. Section 6.2 discusses possible extensions when the empirical histogram deviates from the expected limit shape.
@@
[…]
@@
[…]
@@
$\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ $1$,  $1$,   $-\frac{3}{2}$ $\gamma(1+\gamma)^{4/3}$,   $\frac{1}{\gamma(1+\gamma)^{2/3}}$,  $-\frac{3}{2}\gamma(1+\gamma)^{2}$
@@
$\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ $\gamma(1+\gamma)^{4/3}$,   $\frac{1}{\gamma(1+\gamma)^{2/3}}$,   $-\frac{3}{2}\gamma(1+\gamma)^{2}$ $\frac{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{4/3}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{4/3}}{\tau_{N}^{5/3}\tau_{M}(\tau_{N}-1)^{1/6}(\tau_{M}-1)^{1/6}}$ , $\frac{\tau_{N}^{1/3}(\tau_{N}-1)^{5/6}(\tau_{M}-1)^{5/6}}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2/3}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{2/3}}$ , $-\frac{3}{2}\frac{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{2}}{\tau_{N}^{2}\tau_{M}\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}}$
@@
Visually, the procedure amounts to plotting the functions $\theta\mapsto\widehat{t}_{\pm}(\theta)$ and finding their intersection with the horizontal line $y=\lambda_{i}$. The resulting interval $[\theta_{-},\theta_{+}]$ serves as the confidence interval for the $i$th largest signal strength $\theta_{i}$. Corollary 4.8 ensures that as $N\to\infty$ we have $\mathrm{Prob}(\theta_{i}\in[\theta_{-},\theta_{+}])\to 1-\alpha$.
@@
3.3. Unknown variance
For the CCA setting in Section 2.4 the asymptotics in Theorem 4.6 do not depend on the noise covariance, i.e., the matrices $A$ and $B$ in (2.14). In contrast, for the three other settings, Sections 2.1, 2.2, and 2.3, the scaling depends on the noise variance, denoted by $\sigma^{2}$. Procedure 3.1 assumes $\sigma^{2}=1$. If $\sigma^{2}\neq 1$ but is known, then the entries of the data matrix $\mathbf{A}$ or $X$ should be divided by $\sigma$ to reduce to the baseline case $\sigma^{2}=1$. If $\sigma^{2}$ is unknown, it must first be estimated.
We propose estimating the variance by discarding $25\%$ of the eigenvalues from both the lower and upper ends, and matching sample moments to their theoretical values in order to solve for $\sigma^{2}$.

### v2

@@ 2.4. Canonical correlation analysis
where $A$ and $B$ are $N\times N$ and $M\times M$ matrices, respectively, $I_{N}$ and $I_{M}$ are identity matrices of $N\times N$ and $M\times M$ dimensions, respectively, and $\mathrm{diag}(\sqrt{\theta_{1}},\dots,\sqrt{\theta_{r}})$ is a rectangular matrix with $\sqrt{\theta_{1}},\dots,\sqrt{\theta_{r}}$ on the first $r$ elements of the main diagonal and 0 everywhere else.
Let $\mathbf{x}$ be an (N+M)(N+M)–dimensional Gaussian mean 0 random vector with covariance $\Omega$, and let $\mathbf{u}$ and $\mathbf{v}$ denote its first $N$ and last $M$ coordinates, respectively. The parameters $\theta_{1},\dots,\theta_{r}$ are the squared canonical correlations between $\mathbf{u}$ and $\mathbf{v}$; see Bykhovskaya and Gorin (2024), as well as classical statistics references such as Thompson (1984); Gittins (1985); Anderson (2003); Muirhead (2009) for detailed introductions to canonical correlation analysis (CCA). Algorithmically, $\theta_{i}$ are the largest eigenvalues of the matrix $(\mathbb{E}\mathbf{u}\mathbf{u}^{\mathsf{T}})^{-1}\mathbb{E}\mathbf{u}\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{T}})^{-1}\mathbb{E}\mathbf{v}\mathbf{u}^{\mathsf{T}}$.
Given $S$ independent samples of $\mathbf{x}$, we construct two matrices: the $N\times S$ matrix $\mathbf{U}$ has $S$ samples of $\mathbf{u}$ as columns and the $M\times S$ matrix $\mathbf{V}$ has $S$ samples of $\mathbf{v}$ as columns. The sample squared canonical correlations $\lambda_{1}\geq\lambda_{2}\geq\dots$ are the eigenvalues of the $N\times N$ matrix $(\mathbf{U}\mathbf{U}^{\mathsf{T}})^{-1}\mathbf{U}\mathbf{V}^{\mathsf{T}}(\mathbf{V}\mathbf{V}^{\mathsf{T}})^{-1}\mathbf{V}\mathbf{U}^{\mathsf{T}}$. Our goal is to estimate $\theta_{1},\dots,\theta_{r}$ from these eigenvalues.
@@ 3. Construction of confidence intervals
In this section we present our algorithm for constructing confidence intervals and explain how they can be interpreted and used to distinguish between noise, non-informative signals, and meaningful signals. We begin by introducing the transition process $\mathcal{T}(\Theta)$ and its properties, and then show how to use it to construct confidence intervals. We present two approaches: The first one is based on pre-tabulated quantiles of $\mathcal{T}(\Theta)$. The second one relies on bootstrap-type methodology. The underlying theorems will be presented in Section 4.
@@ 3.1. Transition process
$\Theta$ $\alpha$ .005 .025 .05 .5 .95 .975 .995 -3.0 -3.85 -3.22 -2.89 -0.96 1.32 1.80 2.78 -2.9 -3.85 -3.22 -2.88 -0.95 1.32 1.80 2.79 -2.8 -3.84 -3.20 -2.86 -0.94 1.33 1.81 2.80 -2.7 -3.83 -3.20 -2.86 -0.93 1.35 1.83 2.82 -2.6 -3.82 -3.19 -2.85 -0.92 1.36 1.85 2.83 -2.5 -3.82 -3.18 -2.84 -0.91 1.38 1.86 2.83 -2.4 -3.80 -3.17 -2.83 -0.89 1.38 1.87 2.85 -2.3 -3.80 -3.16 -2.82 -0.89 1.42 1.91 2.90 -2.2 -3.79 -3.15 -2.82 -0.87 1.42 1.90 2.86 -2.1 -3.77 -3.13 -2.79 -0.85 1.44 1.93 2.94 -2.0 -3.75 -3.12 -2.78 -0.84 1.46 1.95 2.97 -1.9 -3.75 -3.11 -2.77 -0.82 1.49 1.98 2.98 -1.8 -3.74 -3.10 -2.75 -0.80 1.52 2.01 3.03 -1.7 -3.73 -3.09 -2.74 -0.78 1.54 2.03 3.05 -1.6 -3.71 -3.06 -2.72 -0.76 1.57 2.07 3.11 -1.5 -3.69 -3.05 -2.70 -0.74 1.61 2.11 3.14 -1.4 -3.67 -3.03 -2.69 -0.71 1.64 2.14 3.19 -1.3 -3.65 -3.01 -2.66 -0.69 1.68 2.19 3.26 -1.2 -3.64 -2.99 -2.65 -0.66 1.72 2.23 3.27 -1.1 -3.61 -2.97 -2.63 -0.63 1.77 2.29 3.37 -1.0 -3.58 -2.94 -2.59 -0.59 1.83 2.35 3.44 -0.9 -3.57 -2.91 -2.56 -0.56 1.88 2.41 3.52 -0.8 -3.55 -2.89 -2.54 -0.53 1.94 2.48 3.62 -0.7 -3.53 -2.87 -2.51 -0.48 2.01 2.57 3.70 -0.6 -3.51 -2.84 -2.49 -0.44 2.08 2.64 3.79 -0.5 -3.49 -2.82 -2.46 -0.39 2.17 2.74 3.93 -0.4 -3.45 -2.77 -2.42 -0.35 2.27 2.85 4.10 -0.3 -3.41 -2.74 -2.38 -0.28 2.38 2.97 4.25 -0.2 -3.37 -2.70 -2.33 -0.22 2.48 3.09 4.39 -0.1 -3.35 -2.66 -2.30 -0.15 2.61 3.24 4.59 0.0 -3.31 -2.62 -2.25 -0.08 2.75 3.40 4.76 0.1 -3.26 -2.57 -2.20 0.00 2.91 3.57 4.96 0.2 -3.22 -2.52 -2.14 0.09 3.08 3.77 5.24 0.3 -3.17 -2.47 -2.09 0.19 3.25 3.95 5.45 0.4 -3.13 -2.40 -2.02 0.29 3.46 4.19 5.72 0.5 -3.08 -2.34 -1.96 0.41 3.67 4.42 5.99 0.6 -3.01 -2.28 -1.89 0.54 3.90 4.68 6.29 0.7 -2.96 -2.21 -1.80 0.69 4.17 4.96 6.60 0.8 -2.90 -2.14 -1.73 0.83 4.41 5.23 6.91 0.9 -2.81 -2.04 -1.62 1.01 4.73 5.57 7.30 1.0 -2.71 -1.95 -1.53 1.18 5.02 5.88 7.62 1.1 -2.63 -1.86 -1.42 1.39 5.37 6.23 8.01 1.2 -2.56 -1.75 -1.31 1.60 5.69 6.59 8.38 1.3 -2.47 -1.64 -1.18 1.84 6.06 6.96 8.80 1.4 -2.37 -1.52 -1.05 2.10 6.44 7.37 9.22 1.5 -2.25 -1.38 -0.89 2.37 6.82 7.79 9.72     $\Theta$ $\alpha$ .005 .025 .05 .5 .95 .975 .995 1.5 -2.25 -1.38 -0.89 2.37 6.82 7.79 9.72 1.6 -2.14 -1.23 -0.73 2.67 7.27 8.22 10.19 1.7 -2.03 -1.08 -0.55 3.00 7.67 8.65 10.66 1.8 -1.88 -0.91 -0.37 3.34 8.12 9.13 11.19 1.9 -1.73 -0.72 -0.14 3.70 8.59 9.62 11.71 2.0 -1.56 -0.53 0.07 4.09 9.09 10.14 12.20 2.1 -1.39 -0.30 0.34 4.50 9.61 10.66 12.84 2.2 -1.21 -0.05 0.62 4.91 10.12 11.20 13.31 2.3 -0.99 0.22 0.92 5.37 10.67 11.75 13.96 2.4 -0.77 0.51 1.24 5.84 11.24 12.36 14.52 2.5 -0.53 0.82 1.59 6.32 11.82 12.94 15.14 2.6 -0.24 1.16 1.96 6.82 12.45 13.57 15.88 2.7 0.04 1.52 2.37 7.35 13.05 14.22 16.52 2.8 0.36 1.91 2.78 7.90 13.68 14.85 17.18 2.9 0.73 2.34 3.23 8.47 14.35 15.55 17.90 3.0 1.08 2.77 3.70 9.04 15.02 16.21 18.60 3.1 1.52 3.27 4.22 9.67 15.71 16.92 19.38 3.2 1.93 3.72 4.72 10.28 16.42 17.63 20.07 3.3 2.41 4.25 5.26 10.93 17.15 18.40 20.85 3.4 2.93 4.79 5.82 11.61 17.91 19.18 21.74 3.5 3.40 5.36 6.41 12.29 18.65 19.93 22.48 3.6 3.96 5.95 7.03 13.01 19.48 20.78 23.37 3.7 4.52 6.59 7.68 13.72 20.26 21.56 24.14 3.8 5.12 7.18 8.30 14.48 21.13 22.47 25.05 3.9 5.76 7.89 9.01 15.25 21.95 23.31 25.98 4.0 6.38 8.57 9.72 16.05 22.83 24.17 26.87 4.1 7.04 9.25 10.41 16.85 23.70 25.06 27.75 4.2 7.70 9.93 11.15 17.68 24.61 25.96 28.68 4.3 8.41 10.69 11.91 18.52 25.55 26.94 29.71 4.4 9.14 11.49 12.73 19.41 26.49 27.88 30.62 4.5 9.93 12.27 13.52 20.28 27.43 28.84 31.60 4.6 10.68 13.09 14.35 21.20 28.43 29.81 32.63 4.7 11.46 13.91 15.19 22.12 29.41 30.84 33.67 4.8 12.23 14.76 16.07 23.06 30.44 31.88 34.68 4.9 13.12 15.63 16.94 24.03 31.52 32.99 35.90 5.0 13.97 16.54 17.87 25.04 32.55 34.02 36.98 5.1 14.88 17.44 18.80 26.04 33.64 35.13 38.05 5.2 15.76 18.39 19.73 27.06 34.74 36.22 39.18 5.3 16.70 19.36 20.71 28.11 35.83 37.34 40.37 5.4 17.67 20.35 21.75 29.19 37.01 38.52 41.57 5.5 18.57 21.35 22.75 30.28 38.14 39.70 42.71 5.6 19.63 22.36 23.78 31.40 39.34 40.91 43.98 5.7 20.67 23.43 24.85 32.54 40.53 42.08 45.15 5.8 21.65 24.46 25.91 33.67 41.72 43.31 46.43 5.9 22.78 25.52 27.00 34.83 42.97 44.54 47.70 6.0 23.82 26.63 28.11 36.03 44.25 45.86 48.95
@@
Table 1. Quantiles of $\mathcal{T}(\Theta)$ for $3\leq\Theta\leq 6$ based on $MC=10^{6}$ Monte Carlo simulations.
$\alpha$ .005 .025 .05 .5 .95 .975 .995
quantile $F_{1}^{-1}(\alpha)$ -4.15 -3.52 -3.18 -1.27 0.98 1.45 2.42
@@
  • For $3\leq\Theta\leq 6$, quantiles are tabulated in Table 1 using the algorithm described in Section 4.1.
@@
    $\mathbb{P}\left\{\mathcal{T}(\Theta)\leq t\right\}\approx\Phi((t-\Theta^{2})/(2\sqrt{\Theta}))\penalty 10000\ .$
@@
    $\mathbb{P}\left\{\mathcal{T}(\Theta)\leq t\right\}\approx F_{1}(t+1/\Theta)\penalty 10000\ ,$
@@
Theorem 4.6 and Corollary 4.8 establish that the approximation (3.1) is valid both when $\theta$ is bounded away from the critical value $\theta^{c}$ and when $\theta$ is close to $\theta^{c}$. One can also show that the approximation remains valid as $\theta\to\infty$ (corresponding to strong signals) [^8] In the case of CCA, $\theta$ represents a correlation and is therefore bounded by $1$.; we omit the proof since the present work focuses on the weak-signal regime. As is evident from the simulations in Figure 1, the quality of the approximation improves as $\theta\to\infty$; we provide one formal result in this direction at the end of Section 9.2.
Our results further show that, in many cases, the approximations for different $\lambda_{i}$ are asymptotically independent. Consequently, we can use (3.1) as a foundation for constructing confidence intervals.
3.2. Confidence intervals with known $\sigma^{2}$: the first algorithm
@@
The first step is to draw a histogram of all eigenvalues $\lambda_{1},\lambda_{2},\dots$. In the settings of (2.1), (2.6), (2.10), or (2.14), the histogram should resemble a known limiting shape; namely, the semicircle law, Marchenko-Pastur law, or Wachter law, depending on the model, as detailed in Table 3, with parameters specified in Table 4, see Appendix 8.4 for more details. If the histogram is reminiscent of one of these shapes, we regard the modelling assumptions as valid and apply Procedure 3.1 to construct confidence intervals. Section 6.2 discusses possible extensions when the empirical histogram deviates from the expected limit shape.
@@
$\frac{1}{2\pi}\sqrt{4-x^{2}}\,\mathbf{1}_{[-2,2]}\,\mathrm{d}x$ $\frac{1}{2\pi}\frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{\gamma^{2}x}\,\mathbf{1}_{[\lambda_{-},\lambda_{+}]}\,\mathrm{d}x$
@@
$\frac{1}{2\pi}\frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{\gamma^{2}x}\,\mathbf{1}_{[\lambda_{-},\lambda_{+}]}\,\mathrm{d}x$ $\frac{\tau_{N}}{2\pi}\frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{x(1-x)}\mathbf{1}_{[\lambda_{-},\lambda_{+}]}\,\mathrm{d}x$
@@
$\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ 11,  11,   $\frac{3}{2}$ $\gamma(1+\gamma)^{4/3}$,   $\frac{1}{\gamma(1+\gamma)^{2/3}}$,  $\frac{3}{2}\gamma(1+\gamma)^{2}$
@@
$\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$ $\gamma(1+\gamma)^{4/3}$,   $\frac{1}{\gamma(1+\gamma)^{2/3}}$,   $\frac{3}{2}\gamma(1+\gamma)^{2}$ $\frac{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{4/3}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{4/3}}{\tau_{N}^{5/3}\tau_{M}(\tau_{N}-1)^{1/6}(\tau_{M}-1)^{1/6}}$ , $\frac{\tau_{N}^{1/3}(\tau_{N}-1)^{5/6}(\tau_{M}-1)^{5/6}}{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2/3}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{2/3}}$ , $\frac{3}{2}\frac{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{2}}{\tau_{N}^{2}\tau_{M}\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}}$
@@
The procedure amounts to plotting the functions $\theta\mapsto\widehat{t}_{\pm}(\theta)$ and finding their intersection with the horizontal line $y=\lambda_{i}$. The resulting $\theta_{-},\theta_{+}]$ serves as the confidence interval for the $i$th signal strength $\theta_{i}$. We have $\mathrm{Prob}(\theta_{i}\in[\theta_{-},\theta_{+}])\to 1-\alpha$ as $N\to\infty$ by Corollary 4.8.
@@
3.3. Bootstrap algorithm
An important and unexpected feature of our asymptotic results (see Theorem 4.6 for details) is that for a fixed index $i$ the asymptotic approximation of $\lambda_{i}$ depends only on $\theta_{i}$, but not on other parameters of the models (2.1), (2.6), (2.10), (2.14), such as $\{\theta_{j}\}_{j\neq i}$ or vectors $\mathbf{u}^{*}_{i}$. Additionally, the rank $r$ does not enter into the formulas. Hence, instead of relying on formulas (3.3) for the functions $\widehat{t}_{\pm}(\theta)$, we can obtain them through a bootstrap procedure using only the $r=1$ case.
Alternative formula-free second step. Consider one of the models (2.1), (2.6), (2.10), or (2.14) with the desired matrix sizes $N$, $S$, and $M$, and instead of the true rank set $r=1$. Fix the direction of the unique signal arbitrarily; for example, set $\mathbf{u}^{*}_{1}$ (and $\mathbf{v}^{*}_{1}$ for the factor model, or the canonical variables for CCA) to the first coordinate vector. The model then depends on a single remaining parameter, $\theta_{1}$. Discretize $\theta_{1}$ on a grid, and for each value compute the largest eigenvalue $\lambda_{1}$ of the corresponding model matrix ($\mathbf{A}$ for the spiked Wigner matrix, $\frac{1}{S}XX^{\mathsf{T}}$ for the spiked covariance and factor models, and $(\mathbf{U}\mathbf{U}^{\mathsf{T}})^{-1}\mathbf{U}\mathbf{V}^{\mathsf{T}}(\mathbf{V}\mathbf{V}^{\mathsf{T}})^{-1}\mathbf{V}\mathbf{U}^{\mathsf{T}}$ for CCA). By repeating sufficiently many Monte-Carlo simulations, compute the $\alpha/2$ and $(1-\alpha/2)$ quantiles of $\lambda_{1}$ for each $\theta_{1}$, which yield the desired $\widehat{t}_{-}(\theta)$ and $\widehat{t}_{+}(\theta)$. This procedure produces the thick gray curves in Figure 1. Then proceed to the third step as in Section 3.2.
An advantage of the algorithm in this section is that (3.3) and Tables 1 and 4 are not required. However, this comes at the cost of running many Monte-Carlo simulations, making the implementation slower than that of the algorithm in Section 3.2.
3.4. Unknown $\sigma^{2}$
For the CCA setting in Section 2.4 the asymptotics in (2.17) does not depend on the noise covariance, i.e., the matrices $A$ and $B$ in (2.14). In contrast, for the other three settings, Sections 2.1, 2.2, and 2.3, the scaling depends on the noise variance, denoted by $\sigma^{2}$. The same holds for Theorem 4.6, which underlies the algorithms for constructing confidence intervals in Sections 3.2 and 3.3. In particular, Procedure 3.1 assumes $\sigma^{2}=1$. If $\sigma^{2}\neq 1$ but is known, then the entries of the data matrix $\mathbf{A}$ or $X$ should be divided by $\sigma$ to reduce to the baseline case $\sigma^{2}=1$. If $\sigma^{2}$ is unknown, it must first be estimated.
We propose estimating the variance by discarding $25\%$ of the eigenvalues at both ends and matching sample moments to their theoretical values to solve for $\sigma^{2}$.